# You Need to Pay Better Attention: Rethinking the Mathematics of Attention Mechanism

### v2

@@ 6 Discussion
There are two limitations in this paper. First, Super Attention currently supports only fixed context length due to the fixed size of $W^{A}$ (see Eq. 22 and LABEL:subfig:_SuperAttention). Nonetheless, these do not affect the advantages of Super Attention in many SotA applications such in ViT. Moreover, this can be addressed using sliding window, which is future works currently in progress. Second, because of limited computational resources, we could only validate our hypotheses on models with up to 60 million parameters trained on datasets with up to 3.6 million examples. Thus, it is crucial to scale the experiments using more computational resources and train large multi-modal and language models using the proposed mechanisms to better understand their effect as the model and dataset sizes scale.
7 Conclusions
We studied the multi-head scaled dot-product attention, demonstrated its inefficiencies, and proposed pathways to mitigate them. Specifically, we introduced three improved SDPAs, listed below, which allow us to reach on-par or better performance compared to standard SDPA with fewer parameters. We validated these on a range of vision and NLP tasks, within our computation budget, with datasets of varying sizes ranging from 50K to 3.6M and models of varying size from 300K to 60M.
Optimised Attention reduces the computational cost of SDPA, without negatively affecting its performance. It achieves this by removing the values’ linear transformation (LABEL:subfig:_OptimisedAttention), reducing SDPA’s size by 25% while performing similarly in vision and natural language tasks.
Efficient Attention reduces the computational cost of SDPA even further (Fig. 1(c)), by also removing the keys’ linear transformation; thus, using 50% less parameters than standard SDPA. Our evaluation indicates that this allows us to reach up to two times compared to standard SDPA, without negatively affecting the performance.
Super Attention improves the performance of SDPA in both vision and NLP tasks by introducing a new linear transformation, which is applied to the values from left (LABEL:subfig:_SuperAttention). This enhances the performance of all previously discussed attention mechanisms; however, we use it on top of Efficient Attention to also benefit from smaller model sizes and faster computation time. Thereby, Super Attention offers superior performance while being smaller than standard SDPA by around 25% and faster by up to 45% depending on the context size and model dimension.
The promising performance of the attention mechanisms introduced here in vision and NLP underscores their versatility and potential to redefine the landscape of AI. In particular, the attention mechanisms introduced here can open new pathways to train and deploy capable models on devices with limited computational resources, such as smartphones and small personal devices and robots.
Acknowledgments and Disclosure of Funding
This work is partially supported by the UK EPSRC via the Centre for Doctoral Training in Intelligent Games and Game Intelligence (IGGI; EP/S022325/1). Some of the experiments in this paper were done with the support of Vultr and Google cloud credits.
Appendix A Societal Impact
This paper has the potential to positively impact society by enabling the training of smaller and more efficient models. The mechanisms introduced in our work are 25% or 50% smaller than the commonly adopted version of the multi-head attention algorithm, require significantly less computation (in terms of FLOPs), and perform on par with or better than the current algorithm. This can help make progress in the democratization of AI by making it more widely accessible to a broader part of society. By reducing the computational resources and energy consumption required to train and deploy state-of-the-art language models and foundational models, our work can help lower the barriers to entry and promote widespread access to these powerful tools.
Furthermore, the increased efficiency of our proposed methods can lead to a reduction in the carbon footprint associated with AI development and deployment, aligning with the growing need for more sustainable and environmentally friendly technologies. As a result, our research has the potential to not only advance the field of AI but also to contribute to a more inclusive and sustainable future for AI-driven technologies.
@@ B.1 Edge Device Performance
Our main motivation for introducing Optimised, Efficient, and Super Attention is to allow running more capable models on edge devices. We calculated the inference times of the Transformer models, we trained before, on a MacBook Pro with an M2 Chip for each task/attention mechanism in Table 4. As expected, Efficient models are the fastest. Also, Super Attention and Optimised Attention models are faster than their standard counterparts with the same number of heads while performing equally well as we discussed before.
@@
1 4.43 34.84 299.26 0.114 1.02
Standard 4 5.27 46.06 323.84 0.183 1.77
8 6.89 (4) 62.08 (4) 341.69 (4) 0.266 (4) 2.84 (4)
1 4.19 33.36 281.14 0.109 1.00
Optimised 4 5.22 44.17 301.30 0.176 1.72
8 6.37 (2) 60.63 (2) 320.49 (3) 0.262 (2) 2.77 (2)
1 3.78 31.50 259.71 0.101 0.93
Efficient 4 4.71 42.16 276.15 0.170 1.66
8 6.10 (1) 58.60 (1) 301.24 (1) 0.256 (1) 2.70 (1)
1 4.21 33.69 264.99 0.112 0.99
Super 4 5.07 44.47 284.49 0.178 1.74
8 6.65 (3) 60.73 (3) 309.72 (2) 0.264 (3) 2.77 (2)
@@ B.2 Speed and Efficiency Comparison
In the main body and other sections of the Appendix, we present comprehensive theoretical comparisons and rigorous experiments on Vision and NLP classification tasks as well as for English-to-Spanish translation to compare the attention algorithms. Optimised Attention and Efficient Attention perform on par with standard attention with 25% and 50% less parameters respectively. In addition, Super Attention outperformed all other algorithms significantly while having 25% fewer parameters compared to standard attention.
As mentioned in the main body, according to the definitions of our proposed algorithms, Efficient, Optimised, and Super Attention mechanisms perform 2,1, and 1 fewer matrix multiplication per head compared to standard attention respectively. Here, we further analyze and compare the required number of FLOPs for completing a single forward and backward pass for all algorithms under study to gain further insight into the efficiency of the proposed algorithms.
@@
As depicted in LABEL:fig:FLOPs_Versus_Proj_Dim, we compare the number of required FLOPs by each attention algorithm when we fixate the sequence length (denoted as $\ell$) and vary the projection dimension. Even though the number of FLOPs scales linearly with the projection dimension for all algorithms, the slope of this increase differs significantly for each algorithm. Specifically, for Efficient Attention, the slope of the line is equal to $9\ell$ while for both Optimised and Super Attention this is equal to $12\ell$ compared to $15\ell$ for standard attention. This means that as we scale the projection dimension the FLOPs required for finishing a forward and backward pass using Efficient Attention increases $\nicefrac{{3}}{{5}}$ as fast as standard attention.
Figure 2: Number of Flops required to complete a single forward plus backward pass for each attention mechanism. While the complexity and therefore, the number of FLOPs increases linearly as the projection dimension increases for all attention mechanisms, the slope of the increase varies significantly as depicted in this plot. Efficient Attention and Super Attention (Optimised Attention is not shown as it is exactly similar to Super Attention) require significantly fewer FLOPs as the projection dimension increases compared to standard attention. Here sequence length is set to 64 ($\ell=64$). Trying different values for $\ell$ changes the scale of the $y$-axis but the chart looks the same.
@@
where $C_{\text{Attn}}$ is the attention algorithm constant which is 15 for standard attention, 12 for Optimised and Super Attention, and 9 for Efficient Attention, and $\ell$, $d_{m}$, and $h$ represent the sequence length, projection dimension, and number of heads consistent with the notation used throughout the paper.
Fig. 4 shows the 3D plot summarising the number of FLOPs for each attention algorithm under varying sequence length and projection dimension in the single head setting. As evident in Fig. 4 and Eq. 24, our proposed algorithms need fewer FLOPs as sequence length increases, which is an important consideration for use in LLMs.
@@
(a) Stn. Attention
(b) Opt. Attention
(c) Eff. Attention
(d) Sup. Attention
Figure 4: 3D plots summarising how many FLOPs are required to complete a forward plus backward pass given the sequence lengths and projection dimensions for each architecture in a single head setting. Efficient Attention followed by Super Attention and Optimised Attention needs significantly fewer FLOPs for completing a forward and backward pass compared to standard attention.
@@ B.3 Vision Transformers
We trained ViT models with different attention mechanisms, all with two attention layers and model dimension $d_{m}=128$. As expected, Super Attention outperforms all other architectures, in terms of accuracy, by at least $2.68\%$ and standard attention by $3.23\%$. The smallest attention layer size belongs to Efficient Attention, which performs on par with standard attention. The complete results are presented in Table 5.
@@
Att. $h$ $d_{m}$ $d_{k}$ Param. Avg. Time (s) Acc. (%) Loss Test Acc. (%) Test. Loss
@@
CIFAR100.
Classifying CIFAR100 images presents considerable difficulty due to the large number of classes in the dataset. This complexity necessitates the maximal utilisation of the attention layers, thereby presenting the perfect challenge for comparing the attention mechanisms discussed here. We trained ViT models with eight attention layers, each with $d_{m}=256$. As presented in Table 1, the Super Attention model surpasses all other architectures achieving $86.51\%$ top-5 accuracy as opposed to standard attention with $80.85\%$ top-5 accuracy. The Efficient Attention model has the smallest attention layer size, only half of that of the standard attention model. For further insight, complete results including analysis of attention algorithms with different heads are presented in Table 6.
Table 6: Averages of different metrics over five runs in the CIFAR100 experiment. The numbers in parentheses indicate the ranking of each mechanism for that metric. An ablation study on the number of heads shows increasing the number of heads helps improve the performance of all algorithms. As expected, the Efficient Attention model has the smallest attention layer size and the Super Attention model performs the best in terms of accuracy and loss.
Att. $h$ $d_{m}$ $d_{k}$ Param. Avg. Time Acc. Loss Top 5 Test Acc. Test Loss Test Top 5
1 256 256 263,168 14.56 38.71 2.35 70.79 33.15 2.87 62.30
2 256 128 263,168 16.22 43.71 2.14 74.98 33.58 2.88 63.18
Stn. 4 256 64 263,168 17.62 47.09 2.00 77.99 34.48 2.91 63.14
8 256 32 263,168 (4) 21.19 (4) 50.38 (2) 1.85 (2) 80.85 (2) 34.64 (3) 2.89 (4) 63.69 (2)
1 256 256 197,376 12.07 39.03 2.35 70.59 31.88 2.90 61.35
2 256 128 197,376 14.70 43.85 2.12 75.82 32.45 2.90 61.15
Opt. 4 256 64 197,376 16.54 48.73 1.91 79.79 35.02 2.87 63.14
8 256 32 197,376 (2) 20.39 (3) 50.19 (3) 1.86 (3) 80.49 (3) 35.03 (2) 2.85 (2) 63.51 (3)
1 256 256 131,584 11.43 40.34 2.28 72.40 31.45 2.89 61.26
2 256 128 131,584 13.68 44.42 2.10 75.92 32.09 2.90 61.24
Eff. 4 256 64 131,584 15.47 46.52 2.01 77.77 33.80 2.89 62.26
8 256 32 131,584 (1) 19.22 (1) 48.96 (4) 1.90 (4) 79.98 (4) 34.54 (4) 2.85 (2) 63.43 (4)
1 256 256 197,890 13.39 50.29 1.84 80.98 35.05 2.85 63.64
2 256 128 197,890 15.32 53.75 1.69 83.21 35.69 2.92 63.55
Sup. 4 256 64 197,890 16.74 55.09 1.64 84.23 35.40 2.88 63.41
8 256 32 197,890 (3) 20.28 (2) 58.68 (1) 1.50 (1) 86.51 (1) 35.97 (1) 2.84 (1) 64.29 (1)
@@
Scaling the vision experiments even further, the ImageNet1k dataset presents much more complexity as the labels comprise 1000 classes. We used a modified ViT-B/16 model architecture, employed different attention mechanisms in its Transformers blocks, and trained the models. Due to our computational constraints, we reduced the number of transformer blocks from 12 to 8, resized the images to $112\!\times\!112$ (instead of the original $224\!\times\!224$) and reduced the patch size from 16 to 8 to enable training on our Nvidia RTX 4090 GPU. Other parameters are similar to the original architecture; specifically, $d_{m}=768$ and $h=12$. Tables 1 and 7 present the results of our experiments on the ImageNet dataset. In both experiments, Super Attention significantly outperforms standard attention (boosting train accuracy 27% and 17.25% in Tables 1 and 7 experiments respectively).
Table 7: Performance of different architectures on the ImageNet dataset. Since different attention layer architectures in the main ImageNet experiment had different numbers of parameters, an interesting ablation study is comparing these architectures when the total number of parameters is very close. To achieve this, we change some hyperparameters like $d_{m}$ or the number of attention layers from the previous experiment. The numbers in parentheses indicate the ranking of each mechanism for that metric. We used a modified ViT-B/16 model, plugged in the attention algorithms in the Transformers block, and trained the models. Super Attention significantly outperforms all other algorithms.
Att. $h$ $d_{m}$ Att. Layers Tot. # Param. Acc. (%) Loss Top 5 Test Acc. (%) Test Loss Test Top 5
@@ B.4.1 Transformer for Text Classification
The IMDB dataset includes 50,000 reviews with binary labels, indicating negative and positive sentiments. The Transformer models, used in this experiment, all have a single attention layer with model dimension and context length 32. The complete results are presented in Table 8.
@@
The Amazon Reviews dataset poses a different challenge than the IMDB dataset as it is a significantly larger dataset with 3,650,000 reviews, containing a wider range of sentiments in $1,2,\dots,5$; higher values indicate more positive sentiment. The Transformer models, used in this experiment, all have three attention layers with model dimension and context length 64. The complete results are presented in Table 9.
@@
Att. $h$ $d_{m}$ $d_{k}$ Param. Avg. Time Acc. Loss Test Acc. Test Loss
1 64 64 16,640 13.81 61.33 0.897 52.84 1.094
Stn. 2 64 32 16,640 16.33 63.61 0.851 52.71 1.091
4 64 16 16,640 (4) 20.38 (4) 62.54 (3) 0.868 (3) 52.74 (4) 1.097 (4)
1 64 64 12,480 12.54 60.71 0.909 52.79 1.093
Opt. 2 64 32 12,480 14.67 62.04 0.884 52.93 1.090
4 64 16 12,480 (2) 19.89 (3) 61.64 (4) 0.876 (4) 52.88 (3) 1.090 (3)
1 64 64 8,320 10.61 62.23 0.873 53.25 1.082
Eff. 2 64 32 8,320 14.05 63.11 0.862 52.67 1.098
4 64 16 8,320 (1) 17.20 (1) 63.55 (2) 0.845 (2) 53.19 (2) 1.080 (2)
1 64 64 12,480 11.96 66.65 0.776 53.87 1.070
Sup. 2 64 32 12,480 15.21 66.30 0.781 54.11 1.064
4 64 16 12,480 (2) 19.77 (2) 66.52 (1) 0.774 (1) 54.25 (1) 1.058 (1)
B.4.2 Transformer for Neural Machine Translation
Table 10: Averages of different metrics over five runs trained on Europarl and Anki English-to-Spanish translation datasets. The numbers in parentheses indicate the ranking of each mechanism for that metric. An ablation study on the number of heads shows increasing the number of heads enhances the performance of all algorithms. Optimised and Efficient Attentions perform on par or better than Standard Attention on most benchmarks with 1/2 and 3/4 as many attention parameters.
Att. $h$ $d_{m}$ $d_{k}$ Param. Avg. Time BLEU Acc. Loss Test BLEU Test Acc. Test Loss

### v3

@@ 6 Discussion and Conclusions
There are two limitations in this paper. First, Super Attention supports fixed context length due to the fixed size of $W^{A}$ (see Eq. 22 and LABEL:subfig:_SuperAttention). Nonetheless, these do not affect the advantages of Super Attention in many SotA applications such as in ViT. Moreover, this can be addressed using a sliding window, which is a future work currently in progress. Second, because of limited computational resources, we could only validate our hypotheses on models with up to 124 million (1.1 billion considering the language model trained in § B.5) parameters trained on datasets with up to 9 billion (30 billion considering § B.5) tokens. Further scaling the experiments beyond our computational resources and training large multi-modal and language models using the proposed mechanisms could facilitate a better understanding of their performance on industrial scales.
Appendix A Reproducibility Statement
The code for all experiments is provided in the supplementary materials. Publicly available datasets are used, with automatic downloads included in the code, except for the Amazon dataset (link in README). The NanoGPT repository (linked in Experimental Setup) details the generative language modelling experiment. Further implementation details are in Section § 4 and § B.3 and B.4.
@@ B.1 Edge Device Performance
Our main motivation for introducing Optimized, Efficient, and Super Attention is to allow running more capable models on edge devices. We calculated the inference times of the Transformer models, we trained before, on a MacBook Pro with an M2 Chip for each task/attention mechanism in Tbl. 5. As expected, Efficient models are the fastest. Also, Super Attention and Optimized Attention models are faster than their standard counterparts with the same number of heads while performing equally well as we discussed before.
@@
1 4.43 34.84 299.26 0.114 0.53
Standard 4 5.27 46.06 323.84 0.183 0.87
8 6.89 (4) 62.08 (4) 341.69 (4) 0.266 (4) 1.34 (4)
1 4.19 33.36 281.14 0.109 0.47
Optimized 4 5.22 44.17 301.30 0.176 0.76
8 6.37 (2) 60.63 (2) 320.49 (3) 0.262 (2) 1.21 (2)
1 3.78 31.50 259.71 0.101 0.44
Efficient 4 4.71 42.16 276.15 0.170 0.72
8 6.10 (1) 58.60 (1) 301.24 (1) 0.256 (1) 1.14 (1)
1 4.21 33.69 264.99 0.112 0.46
Super 4 5.07 44.47 284.49 0.178 0.74
8 6.65 (3) 60.73 (3) 309.72 (2) 0.264 (3) 1.19 (2)
@@ B.2 Speed and Efficiency Comparison
In the main body and other sections of the Appendix, we present comprehensive theoretical comparisons and rigorous experiments on Vision and NLP classification tasks as well as for English-to-Spanish translation to compare the attention algorithms. Optimized Attention and Efficient Attention perform on par with standard attention with 25% and 50% less parameters respectively. In addition, Super Attention outperformed all other algorithms significantly while having 25% fewer parameters compared to standard attention.
As mentioned in the main body, according to the definitions of our proposed algorithms, Efficient, Optimized, and Super Attention mechanisms perform 2,1, and 1 fewer matrix multiplication per head compared to standard attention respectively. Here, we further analyze and compare the required number of FLOPs for completing a single forward and backward pass for all algorithms under study to gain further insight into the efficiency of the proposed algorithms.
@@
As depicted in LABEL:fig:FLOPs_Versus_Proj_Dim, we compare the number of required FLOPs by each attention algorithm when we fixate the sequence length (denoted as $\ell$) and vary the projection dimension. Even though the number of FLOPs scales linearly with the projection dimension for all algorithms, the slope of this increase differs significantly for each algorithm. Specifically, for Efficient Attention, the slope of the line is equal to $9\ell$ while for both Optimized and Super Attention this is equal to $12\ell$ compared to $15\ell$ for standard attention. This means that as we scale the projection dimension the FLOPs required for finishing a forward and backward pass using Efficient Attention increases $\nicefrac{{3}}{{5}}$ as fast as standard attention.
Figure 5: Number of Flops required to complete a single forward plus backward pass for each attention mechanism. While the complexity and therefore, the number of FLOPs increases linearly as the projection dimension increases for all attention mechanisms, the slope of the increase varies significantly as depicted in this plot. Efficient Attention and Super Attention (Optimized Attention is not shown as it is exactly similar to Super Attention) require significantly fewer FLOPs as the projection dimension increases compared to standard attention. Here sequence length is set to 64 ($\ell=64$). Trying different values for $\ell$ changes the scale of the $y$-axis but the chart looks the same.
@@
where $C_{\text{Attn}}$ is the attention algorithm constant which is 15 for standard attention, 12 for Optimized and Super Attention, and 9 for Efficient Attention, and $\ell$, $d_{m}$, and $h$ represent the sequence length, projection dimension, and number of heads consistent with the notation used throughout the paper.
Fig. 2 shows the 3D plot summarizing the number of FLOPs for each attention algorithm under varying sequence length and projection dimension in the single head setting. As evident in Fig. 2 and Eq. 24, our proposed algorithms need fewer FLOPs as sequence length increases, which is an important consideration for use in LLMs.
@@
(a) Standard Att.
(b) Optimized Att.
(c) Efficient Att.
(d) Super Att.
Figure 7: 3D plots visualizing the number of FLOPs for each variant in a forward + backward pass given different sequence lengths and projection dimensions in single-head setting. Efficient Att. followed by Super and Optimized Att. needs substantially fewer FLOPs for completing a forward + backward pass compared to standard attention.
Figure 8: Performance of different architectures on the Amazon Reviews Classification task as the size of the models increases from 5 Million parameters to 25 Million parameters. The results point to the overparameterization of the Standard Attention as it puts an additional computational burden which is not accompanied with better performance in terms of accuracy or loss.
@@ B.3 Vision Transformers
We trained ViT models with different attention mechanisms, all with two attention layers and model dimension $d_{m}=128$. As expected, Super Attention outperforms all other architectures, in terms of accuracy, by at least $2.68\%$ and standard attention by $3.23\%$. The smallest attention layer size belongs to Efficient Attention, which performs on par with standard attention. The complete results are presented in Tbl. 6.
@@
Att. $h$ $d_{m}$ $d_{k}$ Param. Avg. Time (s) Acc. (%) Loss Val Acc. (%) Val Loss
@@
Scaling the vision experiments even further, the ImageNet1k dataset presents much more complexity as the labels comprise 1000 classes. We used a modified ViT-B/16 model architecture, employed different attention mechanisms in its Transformers blocks, and trained the models. Due to our computational constraints, we reduced the number of transformer blocks from 12 to 8, resized the images to $112\!\times\!112$ (instead of the original $224\!\times\!224$) and reduced the patch size from 16 to 8 to enable training on our Nvidia RTX 4090 GPU. Other parameters are similar to the original architecture; specifically, $d_{m}=768$ and $h=12$. Tbls. 4 and 7 present the results of our experiments on the ImageNet dataset.
Table 7: Performance of different architectures on the ImageNet dataset. Since different attention layer architectures in the main ImageNet experiment had different numbers of parameters, an interesting ablation study is comparing these architectures when the total number of parameters is very close. To achieve this, we change some hyperparameters like $d_{m}$ or the number of attention layers from the previous experiment. The numbers in parentheses indicate the ranking of each mechanism for that metric. We used a modified ViT-B/16 model, plugged in the attention algorithms in the Transformers block, and trained the models. Super Attention significantly outperforms all other algorithms. Unlike the results reported in Tbl. 4 in the main body, the models in this ablation experiment are not pre-trained on ImageNet21K (as such the accuracies and validation accuracies are lower compared to the ones with pre-training).
Att. $h$ $d_{m}$ Att. Layers Tot. # Param. Acc. (%) Loss Top 5 Val Acc. (%) Val Loss Val Top 5
@@
Val. results in Tbls. 4, 6 and 7 refer to models’ performances on the official validation set for ImageNet1K, and the official tests sets for MNIST and CIFAR100 datasets.
@@ B.4.1 Transformer for Text Classification
The IMDB dataset includes 50,000 reviews with binary labels, indicating negative and positive sentiments. The Transformer models, used in this experiment, all have a single attention layer with model dimension and context length 32. The complete results are presented in Tbl. 8.
@@
The Amazon Reviews dataset poses a different challenge than the IMDB dataset as it is a significantly larger dataset with 3,650,000 reviews, containing a wider range of sentiments in $1,2,\dots,5$; higher values indicate more positive sentiment. The Transformer models, used in this experiment, all have three attention layers with model dimension and context length 64. The complete results are presented in Tbl. 9.
@@
Att. $h$ $d_{m}$ $d_{k}$ Param. Avg. Time Acc. Loss Val Acc. Val Loss
1 128 128 66,048 42.06 86.76 0.31 62.32 0.87
Stn. 2 128 64 66,048 50.91 87.13 0.30 63.66 0.81
4 128 32 66,048 (4) 66.97 (4) 88.49 (3) 0.25 (3) 65.55 (4) 0.77 (3)
1 128 128 49,536 38.68 89.41 0.25 63.03 0.82
Opt. 2 128 64 49,536 47.97 90.48 0.23 65.98 0.75
4 128 32 49,536 (3) 61.75 (3) 89.56 (1) 0.23 (1) 65.67 (2) 0.75 (2)
1 128 128 33,024 34.82 88.56 0.27 63.42 0.81
Eff. 2 128 64 33,024 42.19 88.36 0.27 63.81 0.80
4 128 32 33,024 (1) 56.44 (1) 86.63 (4) 0.29 (4) 65.58 (3) 0.77 (3)
1 128 128 42,336 37.78 89.11 0.26 65.73 0.74
Sup. 2 128 64 42,336 46.24 88.41 0.27 67.22 0.73
4 128 32 42,336 (2) 59.86 (2) 88.56 (2) 0.24 (2) 68.10 (1) 0.71 (1)
B.4.2 Transformer for Machine Translation
Table 10: Averages of different metrics over five runs trained on Europarl and Anki English-to-Spanish translation datasets. The numbers in parentheses indicate the ranking of each mechanism for that metric. An ablation study on the number of heads shows increasing the number of heads enhances the performance of all algorithms. Optimized and Efficient Attentions perform on par or better than Standard Attention on most benchmarks with 1/2 and 3/4 as many attention parameters.
Att. $h$ $d_{m}$ $d_{k}$ Param. Avg. Time BLEU Acc. Loss Val BLEU Val Acc. Val Loss